# SplAgger: Split Aggregation for Meta-Reinforcement Learning

### 05 Mar 2024 v1

@@ 5.2 Memory Benchmarks
We additionally conduct tests on two environments, T-LS and MC-LS, proposed by Beck et al. (2020). Both of these environments were designed to test long-term memory, and the latter has been used previously in meta-RL (Beck et al., 2023b). In the T-LS environment, the agent inhabits a T-shaped maze, called a T-Maze. The agent is shown a signal and then deterministically stepped down a corridor of length 100. At the end of the corridor, that agent opens one of two doors. If the door matches the signal, then the agent receives a reward of 4, and if not, it receives a penalty of -3. In between the start and end of the corridor, observations are augmented by a random variable $n\in\{-1,1\}$, sampled uniformly at random. The agent performs adapts across four sequential episodes, which together constitute a single meta-episode. The MC-LS environment is designed to challenge an agent’s long-term memory based on visual cues from Minecraft. The task involves navigating a sequence of 16 rooms. In each room, the agent must choose to go either left or right around a central column. This decision is based on the column’s material: diamond or iron. Discrete actions allow for a finite set of observations. When the agent makes a decision in line with the observed material, it earns a reward of 0.1. The final decision at the end of the sequence is dictated by a color signal (red or green) presented at the beginning, which specifies the task. As in the T-Maze, the agent receives a high reward (4) for correct final action and a significant penalty (-3) for an incorrect one. The agent performs adapts across two sequential episodes, which together constitute a single meta-episode.
Results in Figure 5 show that SplAgger achieves the highest sample efficiency on both environments. The RNN and PEARL are not able to learn the optimal policy within the allotted number of frames. The reasons for the failure of the the RNN are discussed in Section 6, while potential reasons for the failure of PEARL are analyzed in the Appendix. While CNP is able to learn optimally, it requires significantly more frames on T-LS. AMRL achieves similar performance. Both AMRL and CNP are similar to our method, SplAgger. However, AMRL differs in its gradient estimation and CNP differs in its use of mean aggregator, instead of max, and its lack of an RNN. To fully understand the contribution of each, we systematically modify each of these components, in isolation, next.
5.3 Alternative Aggregation
In this section we introduce modifications of SplAgger, and three new environments designed to systematically test these components.
Planning Game.
First, we evaluate on Planning Game, which is adapted from Ritter et al. (2021). In this environment, the agent inhabits a $3\times 3$ grid that wraps around each side, and the agent must navigate to a goal. The goal location, observed by the agent, changes multiple times during a single episode. The MDP is not defined by the goal location, shown to the agent, but rather by a changing transition function. The state in each cell of the grid changes in each task. A task is thus defined by a permutation of the 9 states. For instance, while the agent may be in the bottom right of the grid in two different tasks, it will observe a different state there in each task. The states themselves are encoded by one-hot encodings. The agent must explore each state once to learn where each state is on the grid. Once it has seen each state, then it can immediately interpret the given goal instruction. Without knowing the state locations, there is a suboptimal policy that does not require remembering each state, namely, re-exploring the entire grid every time a goal is reached.
In this environment, we evaluate alternative methods for combining the RNN with the order-invariant aggregators. We add to our evaluation AMRL without an RNN (i.e., just a max operator and the ST gradient modification), SplAgger without any RNN (i.e., just a max operator), and SplAgger with an RNN used to encode $e_{t}$ but with out the split aggregation. Note that for models where we remove the RNN, we also remove the split connection, since the split connection only serves a purpose in conjunction with the RNN.
Results in Figure 5(a) show that SplAgger and RNN learn the fastest in this domain. Both AMRL and SplAgger without the split connection, which leaves some neurons aggregated, learn a suboptimal policy. Specifically, these methods learn to re-explore the entire grid every time a goal is reached, which we confirm both qualitatively by examining roll-outs, and quantitatively in Section 6. This demonstrates the detrimental affects of the AMRL gradient modification and the benefit of the split aggregation, which motivates SplAgger. We also see that AMRL without an RNN fails to learn any reasonable policy, achieving near zero reward. Since the only difference between this and SplAgger without an RNN is the use of the ST gradient modification, this shows strong evidence of the detrimental effects of the gradient modification in AMRL. The cause of this effect will be analysed in Section 6. While both SplAgger with an RNN and SplAgger without an RNN outperform the sub-optimal exploration policy eventually, and achieve similar returns ultimately, SplAgger with an RNN learns faster initially. This is surprising, since the optimal policy does not require the permutation variance of the RNN. In this environment, however, the suboptimal exploration policy is easier to learn and does require permutation variance. By examining roll-outs, we verify that SplAgger learns this policy early in training. We will discuss this further in Section 6.
@@ 5.3 Alternative Aggregation
Figure 6: Ablations show SplAgger to achieve returns that are equal to or higher than other methods. The Planning Game shows the importance incorporating RNNs with split aggregation and the failure of gradient modification as in AMRL. The T-Maze Agreement domain shows the max aggregation in SplAgger to be beneficial and SplAgger to also improve the performance of average aggregation. The T-Maze Latent environment shows that SplAgger is able to make the max aggregator robust, even in environments where the average aggregator is superior.
T-Maze Agreement.
Second, we evaluate on the T-Maze Agreement in order to investigate the specific permutation invariant operator, $\bigoplus$. In this environment, the agent receives two binary signals: one at the beginning of the maze and one in the middle of the maze. The agent must open a door depending on whether the signals agree or disagree. While the max aggregation can easily identify the support of the state distribution in the data (Beck et al., 2020), and thus easily identify which signals have been seen in each state, the average aggregation must learn to adjust the representation of all states in order to interpret the average. Thus, we hypothesize that this environment should be relatively easier for aggregation with the max operator and harder for aggregation with the mean operator.
We evaluate the average aggregator, using a mean for the operator, in addition to three novel aggregators. First the avgmax aggregator averages half of the neurons and computes a maximum over the other half. Second, the softmax aggregator computes an average over the inputs where the weights are determined by the softmax of the aggregates themselves. In order to aggregate online, in $O(1)$ memory, we store both the sum of the weighted aggregates, i.e., $n=e_{0}\textrm{exp}(e_{0}/\eta)+...+e_{t}\textrm{exp}(e_{t}/\eta)$, and the sum of the weights seen so far, i.e., $d=\textrm{exp}(e_{0}/\eta)+...+\textrm{exp}(e_{t}/\eta)$, where $\eta$ is a learnable temperature parameter. The output is then the quotient of the weighted aggregates and the weights, $n/d$. This aggregator interpolates between the average and max. The initialization of the temperature is important, as it defines the interpolation, and is set to 0.1. Tuning information is available in the appendix. Finally, we compare to a wsoftmax aggregator. This aggregator is the same as softmax, but the weights are predicted separately from the weighted encodings. Half of the input neurons are aggregated, and the other half are used to compute the logits for the softmax weights. An additional version of this aggregator without the softmax is investigated in the Appendix.
Results in Figure 5(b) show that SplAgger achieves the highest returns, demonstrating the superiority of max aggregation in this environment. While the RNN and average variant of SplAgger are able to learn, they require more frames, as expected. The avgmax, softmax, and wsoftmax all learn almost as quickly as SplAgger. Thus, we see that SplAgger is fairly robust to the choice of operator in this domain, as long as it computes some information about the maximum. Additionally, we see that SplAgger with average aggregation (SplAgger-avg) outperforms using the average operator alone (SplAgger-noRNN-avg), which did not learn the optimal policy in the allotted time, suggesting that SplAgger makes the sequence model more robust to a poor choice in the operator.
T-Maze Latent.
Finally, we evaluate on T-Maze Latent, which also modifies the T-Maze environment in order to investigate the specific permutation invariant operator. In this environment, the agent receives an indicator at every timestep. This indicator is either drawn from ${1,-1}$ with a 50% chance of 1 or a 70% chance of 1, depending on the task. The average aggregator should quickly reveal the latent variable, as the variance of the mean decreases, whereas the max aggregator should have a more difficult time counting the occurrence of indicators in different states. Thus, we hypothesize that this environment should be relatively harder for max aggregation methods and easier for average aggregation.
Results in Figure 5(c) show that while SplAgger without an RNN does have a harder time learning, all of the aggregators with RNNs and split aggregation learn at approximately the same rate. We hypothesize that here, SplAgger is able to fall back upon leveraging the RNN. This demonstrates that the combination of the RNN and split aggregation in SplAgger are remarkably robust to different environments, and justify both our aggregation method and use the max aggregator.
@@ 6 Analysis
In this section we analyze different sequential models to gain insights into their performance. We investigate why RNNs remain useful in some cases, even when permutation invariance should be sufficient, and why AMRL performs poorly in our experiments. Specifically, we find that and we find permutation variance to be useful when there exist permutation variant sub-optimal policies. Additionally, we find that both AMRL and SplAgger prevent certain types of gradient decay, but AMRL also causes other gradients to explode.
@@ 6.1 Learning Suboptimal Policies
Figure 7: Here we show results on a modified Planning Game. If no observation is given to the agent, an RNN is required to learn the proper exploration strategy. This is the same exploration as required by the easier sub-optimal in the original Planning Game.
While sensitivity to permutation is not required to learn optimal policies in meta-RL, we find that the RNN improves sample efficiency on the Planning Game. As discussed in Section 5, the Planning Game has a suboptimal policy that re-explores after every goal is reached. To do this, the agent must remember where it is in a sequence of exploratory actions, and then restart the sequence when a new goal is found. While SplAgger without an RNN can surpass the suboptimal policy eventually, it also achieves larger regret than an RNN, since the RNN learns the suboptimal policy faster. Thus, RNNs can enable lower regret when there is an easier permutation variant suboptimal policy. We hypothesize that the reason SplAgger without an RNN cannot learn the suboptimal policy is that the max aggregation can record that a goal has been reached, but cannot identify when that goal was reached, or that it was just reached.
To confirm this, we perform an additional experiment, in which the agent receives no observation of the state, creating partial observability. All the agent observes is whether it is currently at the goal state. Because the agent does not observe the state, there is no way for it to distinguish which MDP it is in, and the problem can no longer be modelled as a distribution of MDPs. In this case, the Markov property does not hold, and permutation invariant aggregation is no longer sufficient for decision making. Still, the optimal policy in this environment requires the same strategy as the suboptimal policy in the Planning Game. The optimal strategy is to explore every state until the goal is reached. When the agent reaches a goal, the goal location is reset, and the agent must restart exploration. In Figure 7, we see that the RNN is able to learn this policy, while SplAgger without an RNN is not. This shows how, even in a distribution of MDPs, where the Markov property holds in each MDP, permutation variance can increase sample efficiency, due to the presence of non-Markovian policies that are suboptimal, but faster to learn.
@@ 6.2 Gradient Decay
Finally, we investigate the gradients of our sequence models to explain why, in some domains evaluated, SplAgger works, while AMRL (Beck et al., 2020) does not. AMRL demonstrates the benefit of its gradient modification by measuring the average gradient of the memory with respect to the encoding of the initial transition, $||\frac{df_{t}}{d\tau_{0}}||_{2}$. AMRL shows that this quantity decays over time for normal sequence models, but not for AMRL, due to the gradient modification. AMRL methods overwrite this gradient to set it equal to the identity. We depict these gradients at initialization in Figure 7(a). (The authors of AMRL also note that another quantity, the signal-to-noise ratio, can additionally have an impact on performance; however we can only recreate this result if we adjust the bias in all models to zero, and find it to be less predictive than the gradients regardless.)
While AMRL prevents gradient decay, it also causes gradient explosion, with respect to the model parameters. Since the number of inputs grows over time, and the norm of the gradients for each input does not shrink, the gradients with respect to the parameters grows. We depict this phenomenon in Figure 7(b). For an input, we sample noise uniformly over $[-1,1]$, and replicate this sample for every dimension in the input to the sequence model. We evaluate over three model initializations.
From these two gradients, it is not clear why SplAgger performs better than an RNN, so we we propose two alternative metrics for evaluation. In Figure 7(c), we plot the gradient of the inputs, over time ($t$), holding the output time (T) fixed: $||\frac{df_{T}}{d\tau_{t}}||_{2}$. Now, we can see that this value is constant for all models, except for the RNN. For an RNN, this value grows as $t$ approaches $T$, implying that, for a fixed output, the inputs become less sensitive backward in time. In other words, by privileging recent transitions, the gradients are not permutation invariant. The gradients are not equal for all inputs, for a given output. In addition to this metric, we can measure the permutation variance directly. In Figure 7(d), we compute the mean difference between encodings of different permutations of inputs at initialization. We also normalize the encodings, to have unit magnitude first. We see that the RNN is the most permutation variant, and sequence models without any RNN, such as SplAgger without an RNN, are the least. Critically, models, like SplAgger, that are the most performant, are not the most or least permutation variant, but rather models that have some components of each.
@@ 7 Conclusion
In this paper we have shown how permutation invariance can be critical when learning to reinforcement learn. We have, for the first time, confirmed this advantage even without the use of task inference objectives. Surprisingly, we also demonstrate that permutation variance can still be useful, both to learn sub-optimal non-Markovian policies early on, and to make the sequence model more robust to the choice of specific aggregation function. Using these insights, we presented SplAgger, making use of split aggregation in order to achieve the best of both methods. Moreover, we have shown that in several domains, popular existing methods, fail, and discussed reasons for the failure of each. In detail, we analyzed how the gradient modification in AMRL causes gradients with respect to the parameters to explode, and measure different types of gradient decay and permutation variance in RNNs. While recent work has demonstrated the efficacy of general-purpose end-to-end objectives in meta-RL learning (Ni et al., 2022; Beck et al., 2023b), and that this can be contingent on the use of hypernetworks for the policy architecture (Beck et al., 2023b), this paper is the first to show that specialized sequence models can still be critical for efficient meta-learning. We present SplAgger, that integrates the best components of the policy architecture, objective, and sequence model into one unified method. We hope that this paper opens the door for further research into sequence models for meta-RL.

### 08 Mar 2024 v2

@@ 5.2 Memory Benchmarks
We additionally conduct tests on two environments, T-LS and MC-LS, proposed by Beck et al. (2020). Both of these environments were designed to test long-term memory, and the latter has been used previously in meta-RL (Beck et al., 2023b). The T-LS environment is depicted in Figure 2(b). The MC-LS environment is designed to challenge an agent’s long-term memory based on visual cues from Minecraft. Environment details can be found in Appendix H.
Results in Figure 5 show that SplAgger achieves the highest sample efficiency on both environments. The RNN and PEARL are not able to learn the optimal policy within the allotted number of frames. The reasons for the failure of the RNN are discussed in Section 6, while potential reasons for the failure of PEARL are analyzed in Appendix D. While CNP is able to learn optimally, it requires significantly more frames on T-LS. AMRL achieves similar performance. Both AMRL and CNP are similar to our method, SplAgger. However, AMRL differs in its gradient estimation and CNP differs in its use of mean aggregator, instead of max, and its lack of an RNN. To fully understand the contribution of each, we systematically modify these components in isolation, in the next section.
(a) Cheetah-Dir
(b) Walker
5.3 Alternative Aggregation
In this section we introduce modifications of SplAgger, and three new environments designed to test these modifications. Since the existing baselines differ in their choice of aggregation function, $g$, encoding of inputs, $e_{t}$, and use of the ST gradient modification, we systematically test these differences here. The most relevant additional baselines are described below, with the rest detailed in Appendix I.
SplAgger-noSplit removes the split aggregation from SplAgger, but still uses max aggregation and an RNN to encode $e_{t}$. Comparing to this method allows us to validate the use of the split aggregation in SplAgger. SplAgger-noRNN removes the RNN from SplAgger in order to test the effects of removing permutation variant components. Since this obviates the need for the split connection, that component is removed as well. Without these components, this method is equivalent to just computing a maximum over linear encodings of each transition, $\tau_{t}$. AMRL-noRNN removes the RNN and split connection from AMRL. Without these components, this method is equivalent computing a maximum over linear encodings of each transition, $\tau_{t}$, along with the ST gradient modification. SplAgger-avg replaces the max operator in SplAgger with an average, in order to test the effects of alternative permutation invariant operators. SplAgger with other operators (avgmax, softmax, wsoftmax) that interpolate between the average and max are evaluated as well, with details in Appendix I.
Planning Game. First, we evaluate on the Planning Game Ritter et al. (2021), in order to evaluate the need for permutation variant components and how to combine them with permutation invariant components. This environment tests an agents ability to discover and remember multiple pieces of information required for subsequent navigation. The Planning Game is useful for evaluation here due to the existence of both a permutation invariant optimal policy and permutation invariant suboptimal policy. The environment is depicted in Figure 2(a) and detailed in Appendix H.
On this domain, we evaluate methods that modify how the RNN is combined with the permutation invariant aggregation. Results in Figure 5(a) show that SplAgger and RNN learn the fastest in this domain. Both AMRL and SplAgger without the split connection learn a suboptimal policy. This demonstrates the detrimental effects of the AMRL gradient modification and the benefit of the split aggregation, which motivates SplAgger. We also see that AMRL without an RNN fails to learn any reasonable policy, achieving near zero reward. Since the only difference between this and SplAgger without an RNN is the use of the ST gradient modification, this shows strong evidence of the detrimental effects of the gradient modification in AMRL. We analyze the causes in Section 6. While both SplAgger with an RNN and SplAgger without an RNN outperform the sub-optimal exploration policy eventually, and achieve similar returns ultimately, SplAgger with an RNN learns faster initially. We discuss this further in Section 6.
@@ 5.3 Alternative Aggregation
T-Maze Agreement.
Second, we evaluate on T-Maze Agreement in order to investigate the specific permutation invariant operator, $\bigoplus$. In this environment, the agent receives two binary signals: one at the beginning of the maze and one in the middle. The agent must open a door depending on whether the signals agree or disagree. This environment is depicted in Figure 2(c). While the max aggregation can easily identify the support of the state distribution in the data (Beck et al., 2020), and thus easily identify which signals have been seen in each state, the average aggregation must learn to adjust the representation of all states to interpret the average. Thus, we hypothesize that this environment is easier for max aggregation and harder for mean aggregation.
On this domain, we evaluate methods that modify the specific permutation invariant operators. Results in Figure 5(b) show that SplAgger achieves the highest returns, demonstrating the superiority of max aggregation in this environment. While the RNN and average variant of SplAgger are able to learn, they require more frames, as expected. The avgmax, softmax, and wsoftmax all learn almost as quickly as SplAgger. Thus, we see that SplAgger is fairly robust to the choice of operator in this domain, as long as it computes some information about the maximum.
T-Maze Latent.
Finally, we evaluate on T-Maze Latent, which also modifies the T-Maze environment in order to investigate the specific permutation invariant operator. In this environment, the agent receives an indicator at every timestep. This indicator is either drawn from $\{0,1\}$ with a 50% chance of 1 or a 70% chance of 1, depending on the task. This environment is depicted in Figure 2(d). The average aggregator should quickly reveal the latent variable, as the variance of the mean decreases, whereas the max aggregator should have a more difficult time counting the occurrence of indicators in different states. Thus, we hypothesize that this environment should be harder for max aggregation methods and easier for average aggregation.
On this domain, as on T-Maze Agreement, we evaluate methods that modify the specific permutation invariant operators. Results in Figure 5(c) show that, surprisingly, all operators used with SplAgger learn at approximately the same rate. We hypothesize that here, SplAgger is able to fall back upon leveraging the RNN. Since the environment was designed to be more difficult for the max operator, we predict that there may be a difference between the operators when the RNN and skip connection of SplAgger are not used. To test this hypothesis, we conduct two additional experiments. We test both SplAgger without an RNN, but still with the default max operator (SplAgger-noRNN), and SplAgger without an RNN but with the average operator (SplAgger-noRNN-avg). Since the removal of the RNN obviates the need for the split aggregation, the split aggregation is removed as well. We see that, as predicted, the method with the average operator performs better than the method with the max operator, when the RNN and split connection are removed. This demonstrates that the combination of the RNN and split aggregation make SplAgger remarkably robust to different environments, and justifies both our aggregation method and the max operator.
@@
Figure 6: SplAgger achieves returns that are equal to or higher than other methods. The Planning Game shows the importance of incorporating RNNs, either in isolation or with split aggregation, and the failure of gradient modification as in AMRL. Combining the RNN and permutation invariant aggregation, without the split connection (SplAgger-noSplit), decreases performance of each. The T-Maze Agreement domain shows the max operator to be beneficial, enabling performance even greater than the RNN when used with SplAgger. The T-Maze Latent environment shows that SplAgger is able to make the max aggregator performant, even in environments where the computing the average alone is superior. Note that each legend shows the additional methods introduced for that experiment, while the legend at the top shows methods from prior experiments.
@@ 6 Analysis
In this section we analyze different sequential models to gain insights into their performance. We investigate why RNNs remain useful in some cases, even when permutation invariance should be sufficient, and why AMRL performs poorly in our experiments. Specifically, we find permutation variance to be useful when there exist permutation variant suboptimal policies that form a useful stepping stone for learning optimal policies later. Additionally, we find that both AMRL and SplAgger prevent certain types of gradient decay, but AMRL also causes other gradients to explode.
@@ 6.1 Learning Suboptimal Policies
While sensitivity to permutation is not required to learn optimal policies in meta-RL, we find that the RNN surprisingly improves sample efficiency on the Planning Game. As discussed in Section 5, the Planning Game has a permutation variant suboptimal policy. This policy re-explores all states after every goal is reached. To do this, the agent must remember where it is in a sequence of exploratory actions, and then restart the sequence when a new goal is found. We examine rollouts to confirm that SplAgger first learns this suboptimal policy. While SplAgger without an RNN can surpass the suboptimal policy eventually, it also achieves achieves lower returns throughout training compared to an RNN, since the RNN learns the suboptimal policy faster. Thus, RNNs can achieve higher returns sooner, when there is a permutation variant suboptimal policy that can act as a stepping stone for learning the optimal policy. We hypothesize that the reason SplAgger without an RNN cannot learn the suboptimal policy is that the max aggregation can record that a goal has been reached, but cannot identify when that goal was reached, or that it was just reached.
Figure 7: Here we show results on a modified Planning Game. If no observation is given to the agent, an RNN is required to learn the proper exploration strategy. This is the same exploration as required by the easier sub-optimal policy in the original Planning Game.
To confirm this, we perform an additional experiment, in which the agent receives no observation of the state, creating partial observability. All it observes is whether it is currently at the goal state. Hence, there is no way for it to distinguish which MDP it is in, and the problem can no longer be modelled as a distribution of MDPs. In this case, the observations are not Markov and permutation invariant aggregation no longer suffices for decision making. Still, the optimal policy in this environment requires the same strategy as the suboptimal policy in the Planning Game: explore every state until the goal is reached. When the agent reaches a goal, the goal location is reset, and the agent must restart exploration. Figure 7 shows that the RNN is able to learn this policy, while SplAgger without an RNN is not. This shows how, even in a distribution of MDPs, where the Markov property holds in each MDP, permutation variance can improve sample efficiency, due to the presence of non-Markov policies that are suboptimal but faster to learn.
@@ 6.2 Gradient Decay
Finally, we investigate the gradients of our sequence models to explain why, in some domains, SplAgger works, while AMRL (Beck et al., 2020) does not. AMRL demonstrates the benefit of its gradient modification by measuring the average gradient of the memory with respect to the encoding of the initial transition, $||\frac{df_{t}}{d\tau_{0}}||_{2}$. AMRL shows that this quantity decays over time for normal sequence models, but not for AMRL, due to the gradient modification. AMRL methods overwrite this gradient to set it equal to the identity. [^1] Beck et al. (2020) also note that the signal-to-noise ratio can also affect performance. However we can only recreate this result by setting the bias in all models to zero, and find it less predictive than the gradients regardless. We depict these gradients at initialization in Figure 7(a), evaluating over three model initializations.
While AMRL prevents gradient decay, it also causes gradient explosion, with respect to the model parameters, $||\frac{df_{t}}{d\theta}||_{2}$. Since the number of inputs grows over time, and the norm of the gradients for each input does not shrink, the gradients with respect to the parameters grows. We depict this phenomenon in Figure 7(b). For an input, we sample noise uniformly over $[-1,1]$, and replicate this sample for every dimension in the input to the sequence model.
From these two gradients, it is not clear why SplAgger performs better than an RNN, so we propose two alternative metrics for evaluation. In Figure 7(c), we plot the gradient of the inputs, over time ($t$), holding the output time (T) fixed: $||\frac{df_{T}}{d\tau_{t}}||_{2}$. This value is roughly constant for all models, except for the RNN. For an RNN, it grows as $t$ approaches $T$, implying that, for a fixed output, the inputs become less sensitive backward in time. In other words, by privileging recent transitions, the gradients are not permutation invariant. The gradients are not equal for all inputs, for a given output. In addition to this metric, we can measure the permutation variance directly. In Figure 7(d), we compute the mean difference between encodings of different permutations of inputs at initialization. We also normalize the encodings, to have unit magnitude first. We see that the RNN is the most permutation variant, and sequence models without any RNN, such as SplAgger without an RNN, are the least. Critically, models like SplAgger that perform best are not the most or least permutation variant, but rather have some components of each.
@@ 7 Conclusion
In this paper we have shown how permutation invariance can be critical when learning to reinforcement learn. We have, for the first time, confirmed this advantage even without the use of task inference objectives. Surprisingly, we also demonstrate that permutation variance can still be useful, both to learn sub-optimal non-Markovian policies early on, and to make the sequence model more robust to the choice of specific aggregation function. Using these insights, we presented SplAgger, making use of split aggregation to achieve the best of both methods. Moreover, we have shown that in several domains, popular existing methods fail, and discussed reasons for the failure of each. We analyzed how the gradient modification in AMRL causes gradients with respect to the parameters to explode, and measure different types of gradient decay and permutation variance in RNNs.